# Image Inpainting based on Visual-Neural-Inspired Specific Object-of-Interest Imaging Technology

### v2

@@ 6 Conclusion
In conclusion, this work proposes Specific Object-of-Interest Imaging (SIOI) as a novel, bio-inspired pre-processing paradigm for object-centric visual tasks. SIOI operates by first isolating and enhancing the target object, providing a high-fidelity prior that dramatically improves the performance of subsequent vision modules. Extensive experiments confirm that the SIOI module is a powerful, plug-and-play front-end that consistently improves various inpainting models by suppressing background interference and providing a coherent object-centric prior. Its design is grounded in the ”object-first” processing hierarchy of the ventral visual pathway, offering a computational solution to the problem of information redundancy.The framework demonstrates strong robustness across challenging conditions, reflecting the resilience of biological attentional systems. Limitations arise from reliance on initial object generation, which may impair boundary guidance. We envision the SIOI paradigm extending beyond inpainting to benefit a wide array of object-centric applications, including segmentation, detection, and video tracking, wherever suppressing background clutter and focusing computational resources on the target is paramount. Future work will explore task-driven modulation and fMRI validation, advancing toward biologically adaptive imaging and broader applications in neuroimaging and perceptual assistance.

### v3

@@ 6 Conclusion
In conclusion, we propose Specific Object-of-Interest Imaging (SIOI), a bio-inspired preprocessing paradigm for object-centric visual tasks. It operates by isolating and enhancing the target object, providing a high-fidelity prior that enhances subsequent vision modules. Extensive experiments confirm SIOI as a powerful, plug-and-play front-end that improves diverse inpainting models by suppressing background interference and providing a coherent object-centric prior. Grounded in the ventral pathway’s ”object-first” hierarchy, it offers a computational solution to information redundancy. The framework demonstrates robustness under challenging conditions, reflecting the resilience of biological attention. A limitation is its reliance on initial object generation, which can impair boundary guidance. The SIOI paradigm can benefit other object-centric tasks like segmentation, detection, and tracking, where suppressing background clutter is critical. Future work includes task-driven modulation and fMRI validation, advancing toward biologically adaptive imaging for neuroimaging and perceptual assistance.